OpenStudy (anonymous):

Using complete sentences, explain the two ways to simplify 320−−−√ and provide the simplified form.

OpenStudy (akashdeepdeb):

Write it in the form of factors \[\sqrt{320} = \sqrt{2*2*2*2*2*2*5} = \sqrt{(2*2)*(2*2)*(2*2)*5} = 8*\sqrt{5}\] That should be your answer! :) To solve sums like these do it in the above method! :)

OpenStudy (akashdeepdeb):

Or you can write \[\sqrt{320} = \sqrt{64*5} = 8*\sqrt{5}\] [Because 64 square root is 8 so it goes out side the root sign] Understood? :)

OpenStudy (anonymous):

i have to explain in complete sentences

OpenStudy (akashdeepdeb):

Then try writing it in words!! B)

OpenStudy (anonymous):

yes! :) What is the exact value of the expression \[11\sqrt{8}+6\sqrt{12}-5\sqrt{2}\]Show your work.

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (akashdeepdeb):

Okay Now I am going to ask you to do this for me now okay? Try breaking down √8 and √12 in the form of [something*√2] ^_^ TRY IT!! :D

OpenStudy (anonymous):

\[\sqrt{8}= like 3 \sqrt{12}=3.5\]

OpenStudy (akashdeepdeb):

Not quite like I did before group everything in two's Like √8 = √2*2*2 = √4*2 [Now how much is square root of 4? 2!! :D] So √8 = 2√2 Now write √12 in this kind of form! :)

OpenStudy (anonymous):

confused.

OpenStudy (akashdeepdeb):

Do you know what cube-roots and square-roots are?

OpenStudy (anonymous):

sqrt yes

OpenStudy (akashdeepdeb):

What is the square root of 64?

OpenStudy (anonymous):

8

OpenStudy (akashdeepdeb):

And how did you get that?

OpenStudy (anonymous):

8x8

OpenStudy (akashdeepdeb):

Wow you are a model? : O

OpenStudy (anonymous):

yes haha

OpenStudy (akashdeepdeb):

Nice Okay Yes you are right about 8x8 but look carefully at this..